OpenStudy (scarlettfarra2000):

What expression is the GCF of the terms of the polynomial? 16x^3+28x^5y

OpenStudy (dogzcatz):

4x^5y

OpenStudy (scarlettfarra2000):

Thank you

OpenStudy (dogzcatz):

np

jhonyy9 (jhonyy9):

x^5 and y why are common factors GCF mean great common factor or not ?

jhonyy9 (jhonyy9):

16x^3 =4*4x^3 28x^5y =4x^3 *7x^2y what are common factors ?

OpenStudy (mathmale):

Scarlett: Please try again to find the GCF in your first posted problem. 'y' is not part of this GCF.

jhonyy9 (jhonyy9):

Scarlett can you solve it now ?

OpenStudy (scarlettfarra2000):

Thank you all for your help sorry got pulled away before I as all the help yes I solved it.